# PEARL: Self-Evolving Assistant for Time Management with Reinforcement Learning

### v1

@@
Abstract Overlapping calendar invitations force busy professionals to repeatedly decide which meetings to attend, reschedule, or decline. We refer to this preference-driven decision process as calendar conflict resolution. Automating such process is crucial yet challenging. Scheduling logistics drain hours, and human delegation often fails at scale, which motivate we to ask: Can we trust large language model (LLM) or language agent to manager time? To enable systematic study of this question, we introduce CalConflictBench, a benchmark for long-horizon calendar conflict resolution. Conflicts are presented sequentially and agents receive feedback after each round, requiring them to infer and adapt to user preferences progressively. Our experiments show that current LLM agents perform poorly with high error rates, e.g., Qwen-3-30B-Think has 35% average error rate. To address this gap, we propose PEARL, a reinforcement-learning framework that augments language agent with an external memory module and optimized round-wise reward design, enabling agent to progressively infer and adapt to user preferences on-the-fly. Experiments on CalConflictBench shows that PEARL achieves 0.76 error reduction rate, and 55% improvement in average error rate compared to the strongest baseline.   Date: January, 17th, 2026 Contact: bl61@illinois.edu, hengji@illinois.edu Project Webpage: https://bx126.github.io/pearl.github.io
@@ 1 Introduction
Receiving overlapped calendar invitations a common problem in modern workplaces. Consider a CEO of a company or the PI of a research lab: They need to coordinate a large amount of events with different stakeholders every day, but their daily working hours are limited. When multiple events conflict with each other, they must decide which event to attend, which to postpone, and which to decline. We refer to this repeated, preference-driven decision problem as calendar conflict resolution.
Automating calendar conflict resolution is important because it quietly drains one’s time and undermines productivity. Scheduling logistics associated with meetings, e.g., coordinating availability or rescheduling around last-minute conflicts, can easily amount up to hours each week; workplace statistic suggests that 43% of professionals spend at least three hours per week on scheduling meetings (Reclaim.ai, 2024; Calendly, 2024; Microsoft WorkLab, 2025). While in practice these decisions are often delegated to human assistants such as administrative staff (U.S. Bureau of Labor Statistics, 2025), it can easily break down at scale. Not only do human assistants frequently confront high volume of tasks, but also coordinate multiple stakeholders’ schedule in order to reliably resolve scheduling logistics. Furthermore, when a conflict occurs, human assistants have to rely on sparse, incomplete signals from past trajectories about what the delegator values to resolve the conflict. This leads to their internal preference model drifting over time, leading to judgments that are distant from the delegator’s preference. This calls for a reliable agent that can resolve calendar conflicts. Concretely, a reliable calendar conflict resolution agent should: (i) model long-term individual preferences from past decisions, (ii) adapt when preferences evolve with new context and constraints, and (iii) resolve each conflict by explicitly grounding decisions in the inferred user priors.
The explosive growth of LLMs has enabled language agents. Their ability to perceive and reason over complex information show promise as intelligent assistants that automate real-world tasks across different domains such as software development, chart generation, film-making, and travel planning (Wang et al., 2024; Li et al., 2025b; Qian et al., 2025; Li et al., 2025a). Yet it remains unclear whether their performance is trustworthy for calendar conflict resolution, where small mistakes compound and mis-modeled preferences directly translate into costly time allocation errors. This motivates a central question:
@@
To enable a systematic investigation of this problem, we introduce CalConflictBench, a benchmark for evaluating language agents on calendar event conflict resolution. CalConflictBench features synthetic users with diverse organizational roles and year-long calendars populated with carefully designed conflict scenarios. Conflict events are presented sequentially over time, and the agent receives feedback after each decision. This interactive setup closely mirrors real-world calendar management, where agents must infer and adapt to user preferences progressively through repeated interaction, rather than relying on fixed or one-shot instructions. Our empirical results show that current LLMs struggle on this task with high error rates. These failures reveal a fundamental limitation: LLM agents have weak ability to infer, retain, and refine preference-driven decision principles over long horizons.
To address this gap, we propose PEARL (Preference Evolving Agent with Reinforcement Learning), a reinforcement learning framework that trains language agents to infer user preferences online and apply them consistently over long-horizon calendar conflicts. PEARL introduces a structured rollout with a persistent external memory, the Strategy Hub, which stores a set of interpretable decision strategies (preference states) and is iteratively retrieved and updated at each round to capture newly revealed user priorities. To make preference learning explicit and stable, we optimize the agent with a curriculum-based reward , gradually shifting emphasis from preference inference in early rounds to preference-consistent decision making in later rounds. Experiment shows that PEARL achieves 0.76 error reduction rate CalConflictBench, and 55% improvement on average error rate compared to the strongest baseline.
@@
  • Benchmark. We construct CalConflictBench, an evaluation suite with synthetic data generation engine and standardized evaluation protocols to systematically evaluate LLM agents on calendar conflict resolution, and we provide an in-depth analysis of their failure modes.
  • Method. We propose PEARL5), a reinforcement learning framework that enables agents to progressively infer and adapt to user preferences on-the-fly with explicit memory module and carefully designed round-wise rewards, improving average error rate by 55% over the strongest baseline on CalConflictBench.
@@ 2 Task Formulation
In this section, we formally define the proposed calendar conflict resolution task. Appendix C.4 illustrated an example data point.
Task Objective. The task is modeled as a sequential decision process with state transitions. As illustrated in Figure 1, the goal of calendar conflict resolution is to construct a valid calendar for a single user by resolving a sequence of event conflicts over time. At each step $t$, the agent is presented with current calendar state $\mathcal{C}_{t}$, and a set of temporally overlapping events $\mathcal{E}_{t}=\{e_{t}^{1},\ldots,e_{t}^{N_{t}}\}$ and must accept exactly one event $e_{t}^{i}\in\mathcal{E}_{t}$, rejecting all others. The objective is to progressively model user preferences through interaction and contextual signals, producing a final calendar state $\mathcal{C}_{T}$ that aligns with the user’s preferences and decision context.
Agent Action Space. At step $t$, the agent is tasked to assign a binary decision $a_{t}^{i}\in\{0,1\}$ to each event $e_{t}^{i}\in\mathcal{E}_{t}$, where $a_{t}^{i}=1$ denotes acceptance and $a_{t}^{i}=0$ denotes rejection. The action must satisfy the constraint $\sum_{i}a_{t}^{i}=1$.
@@ 3 CalConflictBench
We introduce CalConflictBench to support the evaluation of proposed task. In the benchmark, we present a synthetic data engine (Section 3.1) for generating realistic, role-specific calendars and a comprehensive evaluation protocol (Section 3.2).
@@ 3.1 Synthetic Data Engine
We construct the synthetic data engine to generate data for training and evaluation. We report the detail of data engine design in Appendix C, and we summarize key steps as following.
@@ 3.2 Evaluation Protocol
Our evaluation is designed to assess the preference-evolving capability of LLM agent, which is whether the agent can infer decision-making principles of users over time. Note that the evaluation designed as single-turn format, and each instance contains history context (past rounds information).
@@
  • Optimal Rank Distance (ORD). For rounds with $M\geq 3$, we ask the agent to produce a ranking $\rho_{t}$ over over the $M=|\mathcal{E}_{t}|$ candidate events. Let $e_{t}^{*}$ be the ground-truth accepted event with 0-indexed position $\mathrm{pos}_{t}(e_{t}^{*};\rho_{t})\in\{0,\dots,M\}$. We define the Optimal Rank Distance ($ORD$) as
@@
  • Average Error Rate of $N$ rounds. The mean decision error across all $N$ rounds in a trajectory, capturing overall long-horizon performance.
@@ 11 $25$ $50$ $75$ $104$ 11 $25$ $50$ $75$ $104$
We follow protocol described in Section 3.2. We vary $M\in\{2,3,4,5\}$ and $W\in\{1,5,10,20\}$ to control the combinatorial difficulty and historical context available at each decision round. More details are reported in Appendix D.
@@
Models. We evaluate a diverse set of strong LLMs as agent base models, spanning open-source, reasoning-oriented, and proprietary families. Our open-source models include Qwen3-8B/14B/30B/30B-Think  Yang et al. (2025), OLMo3-7B/OLMo3-32B-Think Olmo et al. (2025), and LLaMA-3.1-8B Grattafiori et al. (2024). We also include GPT5-nano, GPT5 OpenAI (2025), and Gemini-2.5-Flash Comanici et al. (2025) for proprietary model families. On top of these base models, we further evaluate representative agentic rollout style prompting, including ReAct Yao et al. (2023) and Memory-Augmented ReAct Zhu et al. (2025).
@@ 4.2 Results and Analysis
Figure 2: Average Error Rate of Qwen3-8b under different the number of conflicting events per round ($M$) (left), and Error Reduction Rate of Qwen3-8B under different evaluation parameters (right).
@@ 5.1 Rollout Design for Preference Inference
Strategy Hub. Long-horizon preference learning via pure in-context history is challenging: As interactions grow, agents must repeatedly rediscover the same preference cues from a lengthy, noisy transcript, and the resulting preference state remains implicit and hard to reuse or update. To address this, we introduce the Strategy Hub ($\mathcal{S}$) as an external memory module that maintains a fixed-size set of decision strategies. Each strategy encodes an user preference state in natural language (See Appendix E.1 for details). The design of $\mathcal{S}$ explicitly separates preference inference—identifying which strategy types matter and assigning their weights—from preference execution—applying these learned priorities to new conflict contexts. This decomposition compresses preference learning into a compact and interpretable state that can be persistently updated across rounds, avoiding brittle reliance on implicit long-context representations.
@@ 5.2 Reward Modeling for Preference-Evolving
  • Decision Reward. To directly optimize preference aligned correctness, we reward agent make correct decision: $r_{t}^{\text{a}}(x,y)\;=\;\mathbb{I}\!\left[a_{t}=a_{t}^{*}\right]$, where $a_{t}^{*}$ denote the ground-truth round decision (accept / decline for events in $\mathcal{E}_{t}$).
@@
  • Strategy Hub Interaction Reward. To encourage deliberate preference retrieval/refinement rather than purely reactive decisions, we reward rounds where the agent performs a valid StrategyHub interaction ($u_{t}\in\{0,1\}$): $r_{t}^{\text{s}}(x,y)\;=\;u_{t}$.
@@
$\lambda_{t}^{\text{r}}=0.5*i_{t},\qquad\lambda_{t}^{\text{s}}=0.5*(1-i_{t}).$
The shaped per-round reward as
@@ 5.3 Experiment
Results and Analysis. Figure 5 reveals a clear separation in adaptation dynamics. The zero-shot baseline stays nearly flat around a high error band across rounds, indicating that simply conditioning on growing history does not reliably improve preference alignment and can even slightly drift (negative ERR in Table 2). In contrast, PEARL exhibits a monotonic reduction in error as number of rounds $N$ increases, suggesting that it is not merely exploiting longer context, but is learning to update its decision policy across decision rounds.
Table 2 further disentangles the sources of gains. Providing the memory module access alone (Zero-shot + StrategyHub) yields only modest improvement (AER. decreases from $0.45$ to $0.41$; ERR. increases from 0.029-0.029 to $0.048$), implying that having an external memory without learning is insufficient for robust preference-evolving. Supervised training (SFT) improves final-round accuracy (with AER. of $0.27$) but still lags behind PEARL (with AER. of $0.12$) and achieves substantially weaker adaptation (ERR. $0.325$ vs. $0.761$). This gap suggests that imitation-style training learns better static decision heuristics, yet struggles with long-horizon credit assignment and compounding preference-dependent errors across decision rounds over long horizon. Notably, PEARL achieves 55% improvement on AER. compared to the strongest baseline.
Overall, these results highlight that preference-evolving behavior requires long-horizon optimization over multi-round trajectories: PEARL can translates the history of previous rounds into measurable error reduction, validating the necessity of reinforcement learning for preference adaptation rather than one-shot prompting or purely SFT.
@@
6 Related Works
LLM-based agents have been developed as intelligent assistants for tool-augmented QA, web browsing, and task-oriented dialog (Wang et al., 2024; Li et al., 2025b; Qian et al., 2025), with frameworks such as ReAct and AutoGPT enabling autonomous behavior by interleaving reasoning and tool use (Yao et al., 2023; Yang et al., 2023). Yet personal time management remains less explored: earlier systems (e.g., Calendar.help) depended on predefined workflows with human-in-the-loop execution (Cranshaw et al., 2017), recent studies begin to investigate LLM-based scheduling agents (Shen et al., 2024; Wijerathne et al., 2025). Our work extends this line to long-horizon calendar conflict resolution where agents must adapt to user-specific preferences over many decisions. Preference alignment is commonly achieved via RLHF, which fine-tunes models using human feedback (Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022); to reduce labeling cost, methods leverage AI-generated principles (e.g., Constitutional AI) (Bai et al., 2022), and self-evaluation/self-correction (Wu et al., 2025). Distinctly, we target sequential preference alignment under long horizons. Since long-horizon learning is hindered by limited context and state retention, prior work explores curriculum learning (Narvekar et al., 2020) and external memory/state tracking (Yan et al., 2025); we design external memory module to accumulate past decisions for preference inference and reuse across rounds.

### v2

@@
Abstract Overlapping calendar invitations force busy professionals to repeatedly decide which meetings to attend, reschedule, or decline. We refer to this preference-driven decision process as calendar conflict resolution. Automating this decision process is crucial yet challenging. Scheduling logistics can drain hours, and human delegation often fails at scale, which motivates us to ask: Can we trust large language models (LLMs) or language agents to manage time? To enable a systematic study of this question, we introduce CalConflictBench, a benchmark for long-horizon calendar conflict resolution. In CalConflictBench, conflicts are presented to agents round-by-round over a calendar year, requiring them to infer and adapt to user preferences progressively. Our experiments show that current LLM agents perform poorly with high error rates, e.g., Qwen-3-30B-Think has an average error rate of 35%. To address this gap, we propose PEARL, a reinforcement-learning framework that (i) augments the language agent with an external preference memory that stores and updates inferred strategies (e.g., attendee priorities, topic importance, time/location preferences), and (ii) optimizes the agent with round-wise rewards that directly supervise decision correctness, ranking quality, and memory usage across rounds. Experiments on CalConflictBench show that PEARL achieves an error reduction rate of 0.76 and a 55% improvement in average error rate compared to the strongest baseline.   Date: January, 17th, 2026 Contact: bl61@illinois.edu, hengji@illinois.edu Project Webpage: https://bx126.github.io/pearl.github.io
@@ 1 Introduction
Receiving overlapping calendar invitations is common in modern workplaces. Consider a CEO of a company or a PI of a research lab, they need to coordinate a large amount of events with different stakeholders every day, but their daily working hours are limited. When multiple events conflict with each other, they must decide which event to attend, which to postpone, and which to decline. We refer to this repeated, preference-driven decision problem as calendar conflict resolution.
Automating calendar conflict resolution is important because it quietly drains time and undermines productivity. Scheduling logistics associated with meetings, e.g., coordinating availability or rescheduling around last-minute conflicts, can easily amount to hours each week. Workplace statistics suggest that 43% of professionals spend at least three hours per week on scheduling meetings (Reclaim.ai, 2024; Calendly, 2024; Microsoft WorkLab, 2025). While in practice these decisions are often delegated to human assistants such as administrative staff (U.S. Bureau of Labor Statistics, 2025), it can easily break down at scale. Not only do human assistants frequently confront a high volume of tasks, but they also must coordinate multiple stakeholders’ schedule in order to reliably resolve scheduling logistics. Furthermore, when a conflict occurs, human assistants have to rely on sparse, incomplete signals about what the delegator values to resolve the conflict. This causes their internal preference model to drift over time, leading to judgments that are distant from the delegator’s preferences. This calls for a reliable agent that can resolve calendar conflicts. Concretely, a reliable calendar conflict resolution agent should: (i) model long-term individual preferences from past decisions, (ii) adapt when preferences evolve with new context and constraints, and (iii) resolve each conflict by explicitly grounding decisions in the inferred user priors.
The explosive growth of LLMs has enabled the development of language agents. Their ability to perceive and reason over complex information shows promise as intelligent assistants that automate real-world tasks across different domains, such as software development, chart generation, film-making, and travel planning (Wang et al., 2024; Li et al., 2025b; Qian et al., 2025; Li et al., 2025a). Yet it remains unclear whether their performance is trustworthy for calendar conflict resolution, where small mistakes compound and mis-modeled preferences directly translate into costly time allocation errors. This motivates a central question:
@@
To enable a systematic investigation of this problem, we introduce CalConflictBench, a benchmark for evaluating language agents on calendar event conflict resolution. CalConflictBench features synthetic users with diverse organizational roles and year-long calendars populated with carefully designed conflict scenarios. Conflict events are presented sequentially over time, and the agent receives feedback after each decision. This interactive setup closely mirrors real-world calendar management, where agents must infer and adapt to user preferences progressively through repeated interaction, rather than relying on fixed or one-shot instructions. Our empirical results show that current LLMs struggle on this task with high error rates. These failures reveal a fundamental limitation: LLM agents have a weak ability to infer, retain, and refine preference-driven decision principles over long horizons.
To address this gap, we propose PEARL (Preference Evolving Agent with Reinforcement Learning), a reinforcement learning framework that trains language agents to infer user preferences online and apply them consistently over long-horizon calendar conflicts. PEARL introduces a structured rollout with a persistent external memory, the Strategy Hub, which stores a set of interpretable decision strategies (preference states) and is iteratively retrieved and updated at each round to capture newly revealed user priorities. To make preference learning explicit and stable, we optimize the agent with a curriculum-based reward, gradually shifting emphasis from preference inference in early rounds to preference-consistent decision making in later rounds. Experiment shows that PEARL achieves an 0.76 error reduction rate on CalConflictBench, and 55% improvement in average error rate compared to the strongest baseline.
@@
  • Benchmark. We construct CalConflictBench, an evaluation suite with a synthetic data generation engine and standardized evaluation protocols to systematically evaluate LLM agents on calendar conflict resolution, and we provide an in-depth analysis of their failure modes.
  • Method. We propose PEARL5), a reinforcement learning framework that enables agents to progressively infer and adapt to user preferences on-the-fly with an explicit memory module and carefully designed round-wise rewards, improving average error rate by 55% over the strongest baseline on CalConflictBench.
@@ 2 Task Formulation
In this section, we formally define the proposed calendar conflict resolution task. Appendix C.4 illustrates an example data point.
Task Objective. The task is modeled as a sequential decision process with state transitions. As illustrated in Figure 1, the goal of calendar conflict resolution is to construct a valid calendar for a single user by resolving a sequence of event conflicts over time. At each step $t$, the agent is presented with the current calendar state $\mathcal{C}_{t}$, and a set of temporally overlapping events $\mathcal{E}_{t}=\{e_{t}^{1},\ldots,e_{t}^{N_{t}}\}$ and must accept exactly one event $e_{t}^{i}\in\mathcal{E}_{t}$, rejecting all others. The objective is to progressively model user preferences through interaction and contextual signals, producing a final calendar state $\mathcal{C}_{T}$ that aligns with the user’s preferences and decision context.
Agent Action Space. At step $t$, the agent is tasked with assigning a binary decision $a_{t}^{i}\in\{0,1\}$ to each event $e_{t}^{i}\in\mathcal{E}_{t}$, where $a_{t}^{i}=1$ denotes acceptance and $a_{t}^{i}=0$ denotes rejection. The action must satisfy the constraint $\sum_{i}a_{t}^{i}=1$.
@@ 3 CalConflictBench
We introduce CalConflictBench to support the evaluation of the proposed task. In the benchmark, we present a synthetic data engine (Section 3.1) for generating realistic, role-specific calendars and a comprehensive evaluation protocol (Section 3.2).
@@ 3.1 Synthetic Data Engine
We construct the synthetic data engine to generate data for training and evaluation. We report the details of data engine design in Appendix C, and we summarize key steps as follows.
@@ 3.2 Evaluation Protocol
Our evaluation is designed to assess the preference-evolving capability of LLM agents, which is whether the agent can infer decision-making principles of users over time. Note that the evaluation is designed in a single-turn format, and each instance contains history context (past-round information).
@@
  • Optimal Rank Distance (ORD). For rounds with $M\geq 3$, we ask the agent to produce a ranking $\rho_{t}$ over the $M=|\mathcal{E}_{t}|$ candidate events. Let $e_{t}^{*}$ be the ground-truth accepted event with 0-indexed position $\mathrm{pos}_{t}(e_{t}^{*};\rho_{t})\in\{0,\dots,M-1\}$. We define the Optimal Rank Distance ($ORD$) as
@@
  • Average Error Rate over $N$ rounds. The mean decision error across all $N$ rounds in a trajectory, capturing overall long-horizon performance.
@@ 11 $25$ $50$ $75$ $104$ 11 $25$ $50$ $75$ $104$
We follow the protocol described in Section 3.2. We vary $M\in\{2,3,4,5\}$ and $W\in\{1,5,10,20\}$ to control the combinatorial difficulty and historical context available at each decision round. More details are reported in Appendix D.
@@
Models. We evaluate a diverse set of strong LLMs as agent base models, spanning open-source, reasoning-oriented, and proprietary families. Our open-source models include Qwen3-8B/14B/30B/30B-Think  Yang et al. (2025), OLMo3-7B/OLMo3-32B-Think Olmo et al. (2025), and LLaMA-3.1-8B Grattafiori et al. (2024). We also include GPT5-nano, GPT5 OpenAI (2025), and Gemini-2.5-Flash Comanici et al. (2025) for proprietary model families. On top of these base models, we further evaluate representative agentic rollout style prompting, including ReAct Yao et al. (2023b) and Memory-Augmented ReAct Zhu et al. (2025).
@@ 4.2 Results and Analysis
Figure 2: Average Error Rate of Qwen3-8b under different numbers of conflicting events per round ($M$) (left), and Error Reduction Rate of Qwen3-8B under different evaluation parameters (right).
@@ 5.1 Rollout Design for Preference Inference
Strategy Hub. Long-horizon preference learning via pure in-context history is challenging: As interactions grow, agents must repeatedly rediscover the same preference cues from a lengthy, noisy transcript, and the resulting preference state remains implicit and hard to reuse or update. To address this, we introduce the Strategy Hub ($\mathcal{S}$) as an external memory module that maintains a fixed-size set of decision strategies. Each strategy encodes a user preference state in natural language (See Appendix E.1 for details). The design of $\mathcal{S}$ explicitly separates preference inference—identifying which strategy types matter and assigning their weights—from preference execution—applying these learned priorities to new conflict contexts. This decomposition compresses preference learning into a compact and interpretable state that can be persistently updated across rounds, avoiding brittle reliance on implicit long-context representations.
@@ 5.2 Reward Modeling for Preference-Evolving
  • Decision Reward. To directly optimize preference-aligned correctness, we reward the agent for making correct decision: $r_{t}^{\text{a}}(x,y)\;=\;\mathbb{I}\!\left[a_{t}=a_{t}^{*}\right]$, where $a_{t}^{*}$ denotes the ground-truth round decision (accept / decline for events in $\mathcal{E}_{t}$).
@@
  • Strategy Hub Interaction Reward. To encourage deliberate preference retrieval/refinement rather than purely reactive decisions, we reward rounds where the agent performs a valid StrategyHub interaction ($u_{t}\in\{0,1\}$): $r_{t}^{\text{i}}(x,y)\;=\;u_{t}$.
@@
$\lambda_{t}^{\text{r}}=0.5*i_{t},\qquad\lambda_{t}^{\text{i}}=0.5*(1-i_{t}).$
The shaped per-round reward is
@@ 5.3 Experiment
Results and Analysis. Figure 5 reveals a clear separation in adaptation dynamics. The zero-shot baseline stays nearly flat around a high error band across rounds, indicating that simply conditioning on growing history does not reliably improve preference alignment and can even slightly drift (negative ERR in Table 2). In contrast, PEARL exhibits a monotonic reduction in error as the number of rounds $N$ increases, suggesting that it is not merely exploiting longer context, but is learning to update its decision policy across decision rounds.
Table 2 further disentangles the sources of gains. Providing the memory module access alone (Zero-shot + StrategyHub) yields only modest improvement (AER. decreases from $0.45$ to $0.41$; ERR. increases from 0.029-0.029 to $0.048$), implying that having an external memory without learning is insufficient for robust preference-evolving. Supervised training (SFT) improves final-round accuracy (with AER. of $0.27$) but still lags behind PEARL (with AER. of $0.12$) and achieves substantially weaker adaptation (ERR. $0.325$ vs. $0.761$). This gap suggests that imitation-style training learns better static decision heuristics, yet struggles with long-horizon credit assignment and compounding preference-dependent errors across decision rounds over long horizon. Notably, PEARL achieves a 55% improvement in AER. compared to the strongest baseline.
Overall, these results highlight that preference-evolving behavior requires long-horizon optimization over multi-round trajectories: PEARL can translate the history of previous rounds into measurable error reduction, validating the necessity of reinforcement learning for preference adaptation rather than one-shot prompting or purely SFT.
@@
6 Related Work
LLM-based agents have been developed as intelligent assistants for tool-augmented question answering, web browsing, and real-world downstream tasks such as recipe generation and profile writing (Li et al., 2025b; Qian et al., 2025; Li et al., 2024).Frameworks such as ReAct and AutoGPT enable autonomous behavior by interleaving reasoning and tool use (Yao et al., 2023b; Yang et al., 2023). Beyond tool-use, recent work casts LLM inference as an explicit planning/search problem, ranging from tree-based deliberation (Yao et al., 2023a) and efficiency-oriented search (Katz et al., 2024) to interactive, code-augmented planners that execute and revise programs as plans (Liu et al., 2025). Complementary approaches learn planning-based reasoning by collecting trajectories and synthesizing process rewards for preference-based training (Jiao et al., 2024). Yet personal time management remains less explored: earlier systems (e.g., Calendar.help) depended on predefined workflows with human-in-the-loop execution (Cranshaw et al., 2017); recent studies begin to investigate LLM-based scheduling agents (Shen et al., 2024; Wijerathne et al., 2025). Our work extends this line to long-horizon calendar conflict resolution where agents must adapt to user-specific preferences over many decisions. Preference alignment is commonly achieved via RLHF, which fine-tunes models using human feedback (Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022); to reduce labeling cost, methods leverage AI-generated principles (e.g., Constitutional AI) (Bai et al., 2022), and self-evaluation/self-correction (Wu et al., 2025). Distinctly, we target preference alignment at test time under long horizons. Since long-horizon learning is hindered by limited context and state retention, prior work explores curriculum learning (Narvekar et al., 2020) and external memory/state tracking (Yan et al., 2025); we design external memory module to accumulate past decisions for preference inference and reuse across rounds.